OpenStudy (anonymous):

What is the radius of a circle with circumference 13ft

hero (hero):

\(C = 2 \pi r\) \(C = 13\) \(13 = 2 \pi r\) \(\dfrac{13}{2\pi} = r\)

OpenStudy (anonymous):

If circumference is 2πr², then radius is r/2π. So radius=13/2π=~20.42